# The thermodynamic uncertainty relation of a quantum-mechanically coupled two-qubit system

### v1

@@ IV Discussions
From the 2D map of $\mathcal{Q}(\delta\omega_{1},\delta\omega_{2})$ (Fig. 3A) and $\mathcal{Q}(\delta\omega_{1})$ along the resonance condition (Fig. 3B) for select values of $\tilde{\Omega}$ and $K$, we find that $\mathcal{Q}$ is minimized along the resonance condition, which reads $\delta\omega_{1}=0$ for small $K(\ll 1)$ and $\delta\omega_{1}\delta\omega_{2}=K^{2}$ for large $K(\gtrsim 1)$. At least two distinct values of $\mathcal{Q}_{\rm min}=1.38$ and 1.25, which are smaller than the classical counterpart ($\mathcal{Q}_{\rm min}=2$), are highlighted in Fig. 3B. To identify the value of $\mathcal{Q}_{\rm min}$ along the resonance condition more systematically, we calculate the diagram of $\mathcal{Q}_{\rm min}$ as a function of $\tilde{\Omega}$ and $K$ by simultaneously scanning $\delta\omega_{1}$ and $\mathcal{A}$ in Fig. 4.
@@
Note that $\langle j_{1}\rangle$ is dictated by the typical current-coherence relation (Eq. (21)), whereas $\langle j_{2}\rangle$ is associated only with the imaginary part of the correlated coherence, ${\rm Im}[\rho^{\rm ss}_{10,01}]$. It is noteworthy that another correlated coherence ${\rm Im}[\rho^{\rm ss}_{11,00}]$ (Fig. S1C) makes no contribution to the current. This is most likely associated with the fact that $\langle j_{2}\rangle>0$ (Fig. 2D) for all $\delta\omega_{1}$ and $\delta\omega_{2}$ is more consistent with ${\rm Im}[\rho^{\rm ss}_{10,01}]$ (Fig. 2C) than ${\rm Im}[\rho^{\rm ss}_{11,00}]$, the latter of which is not always positive (Fig. S1C). Despite the inherent asymmetry, the currents generated from the 1st and 2nd qubits become comparable when the two qubits are strongly coupled at large $K$ (data points in green, Fig. 5B).
@@ IV.3 TUR of subsystem and comparison of TQS with two coupled noisy oscillators
(i) For $K_{os}=0$, the two oscillators exhibit phase dynamics fully decoupled from each other [15]. As a result, the total entropy production, the variance and the mean of the phase angle for $\tau\gg 1$ from each oscillator ($i=1$ and 2) are given by
@@
Thus, the uncertainty product for each oscillator is
@@
To draw an analogy with the TQS where only the 1st qubit is irradiated by the external field, we set $\Omega_{1}$ finite and $\Omega_{2}\rightarrow 0$ for the two coupled oscillators. Two limiting cases of the oscillators and their comparisons with the TQS illuminate the fundamental differences between the couplings in the TQS and in the classical oscillators. (i) For $K_{os}\rightarrow 0$, the uncertainty product of the 1st oscillator is minimized to the TUR bound, whereas the 2nd oscillator with $\Omega_{2}\rightarrow 0$ is merely subject to the thermal noise, so that we get $\mathcal{Q}_{1}=2$ and $\mathcal{Q}_{2}\gg 1$. This trend is also observed in the TQS when $\tilde{\Omega}$ is finite and $K\rightarrow 0$. Figure 6A shows the uncertainty products of the total system ($\mathcal{Q}$) and the subsystems ($\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$) as a function of the detunings. As previously discussed, the TQS effectively behaves as two non-interacting qubits. The 1st qubit subject to the external field acquires the minimal uncertainty product $\mathcal{Q}_{1}\approx 1.25$ at $\delta\omega_{1}\approx 0$. Meanwhile, as the 2nd qubit is effectively decoupled from the 1st qubit with no source of driving, its uncertainty product is large ($\mathcal{Q}_{2}\gg 1$, Fig. 6A). (ii) For $K_{os}\gg|\Omega_{1}-\Omega_{2}|\approx\Omega_{1}$, the two oscillators are fully synchronized. Then, the current, its fluctuations, and the entropy production are reduced to half the values, yielding the uncertainty product minimized to its bound, $\mathcal{Q}_{1}=\mathcal{Q}_{2}=2$. Interestingly, in the corresponding regime for TQS ($K\gg 1$), Fig 6B indicates that the uncertainty product of each qubit is significantly large ($\mathcal{Q}_{1,2}\gg 1$) particularly under off-resonance condition. Even under the resonance condition, each qubit’s uncertainty product always exceeds 2, while the uncertainty product for the total system is still around the minimal bound ($\mathcal{Q}\approx 1.25$) (see also Fig. S4 and Fig. 2D, which give rise to $\mathcal{Q}$ and $\mathcal{Q}_{1,2}$). Note that the correlated coherence, particularly the superposition of |$\ket{01}$ and |$\ket{10}$ basis states (${\rm Im}[\rho_{10,01}]$, see Fig. 2C) contributes to the current from each qubit (Eqs. 31 and  32), while the total current originates only from the coherences of the 1st qubit (Eq. (30)). This is a feature unique to the TQS, but lacking in the two coupled oscillators.
@@ IV.4 Effect of entanglement on the TUR bound
Any meaningful signature of entanglement is found for $K\geq 1$ and $\tilde{\Omega}\geq 1$ (Fig. 7). The parameter space giving rise to a strong entanglement does not necessarily coincide with the region with small $\mathcal{Q}$ (compare Fig. 3A and Fig. 7). As vividly shown by the map of the smallest eigenvalue of the partially transposed density matrix ($\Lambda(\delta\omega_{1},\delta\omega_{2})$), the entanglement between the two qubits characterized with $\Lambda<0$ is the most apparent at $\tilde{\Omega}=10$ and $K=10$; yet, it contributes little to loosening the TUR bound. The overlap between the two maps simultaneously satisfying $\mathcal{Q}(\delta\omega_{1},\delta\omega_{2})\lesssim 2$ and $\Lambda(\delta\omega_{1},\delta\omega_{2})<0$, quantified using the overlap parameter $\chi=\Theta(2-\mathcal{Q})(2/\mathcal{Q})\Lambda$, is shaped along the resonance condition ($\delta\omega_{1}\delta\omega_{2}=K^{2}$) only for $\tilde{\Omega}=1$ and $K=1$, 10 (Fig. 7B).
@@ V Summary
    Provided that the external field is strong enough ($\tilde{\Omega}\geq 1$), a strong coupling ($K\gg 1$) between the two qubits is essential for creating an entanglement in the TQS, and for transmitting excitation and coherence from the 1st qubit to the 2nd qubit (compare Fig. 2A, B with Fig. S1 A, B). Strong coupling, however, attenuates the current from the 1st qubit while increasing the photon current from the 2nd qubit (Fig. 2D).
@@
    The most effective suppression of photon current fluctuations, which gives rise to TUR violation, is achieved when $\tilde{\Omega}$ is moderate, and the system satisfies the resonance condition ($\delta\omega_{1}\delta\omega_{2}=K^{2}$).
@@
Here are our final remarks. It is expected that the suboptimal bound of the uncertainty product, $\mathcal{Q}_{K\gg 1,\tilde{\Omega}>3,{\rm min}}^{\rm TQS}\approx 1.36$, is specific to the TQS. Although Eq. (44) is obtained for TQS, we put forward that the relation is general enough to be extended to multi-qubit systems. The TLS sets the fundamental limit of TUR for multi-qubit systems.
@@ VI.2 Method of cumulant generating function
where we have introduced the total emission and absorption operator from the two qubits, ±=±$\mathcal{L}_{\pm}=\mathcal{L}^{1}_{\pm}+\mathcal{L}^{2}_{\pm}$ and defined the remaining elements as $\mathcal{L}_{0}\equiv\mathcal{L}-\mathcal{L}_{+}-\mathcal{L}_{-}$. The two last terms on the right hand side of Eq. (49) express the contributions of emission and absorption to the evolution of $\tilde{\rho}(n,t)$ with $\mathcal{L}_{\pm}$. For the TQS, these operators can be written explicitly as:
@@
To obtain the current from each qubit, the total absorption and emission operator can be redefined as ±=±$\mathcal{L}_{\pm}=\mathcal{L}^{1}_{\pm}$ or ±$\mathcal{L}^{2}_{\pm}$ instead of summing up the entire contribution from the two qubits as ±=±$\mathcal{L}_{\pm}=\mathcal{L}^{1}_{\pm}+\mathcal{L}^{2}_{\pm}$. The expressions for the Liouville operators associated with emissions ($\mathcal{L}_{+}^{1}$, $\mathcal{L}_{+}^{2}$) and absorptions ($\mathcal{L}_{-}^{1}$, $\mathcal{L}_{-}^{2}$) given in Eq. (50) offer the expressions in Eqs. (28) and (29).

### v2

@@ IV Discussions
From the 2D map of $\mathcal{Q}(\delta\omega_{1},\delta\omega_{2})$ (Fig. 3A) and $\mathcal{Q}(\delta\omega_{1})$ along the resonance condition (Fig. 3B) for select values of $\tilde{\Omega}$ and $K$, we find that $\mathcal{Q}$ is minimized along the resonance condition, which reads $\delta\omega_{1}=0$ for small $K(\ll 1)$ and $\delta\omega_{1}\delta\omega_{2}=K^{2}$ for large $K(\gtrsim 1)$. At least two distinct values of $\mathcal{Q}_{\rm min}=1.38$ and 1.25, which are smaller than the classical counterpart ($\mathcal{Q}_{\rm min}=2$), are highlighted in Fig. 3B. To identify the value of $\mathcal{Q}_{\rm min}$ along the resonance condition more systematically, we calculate the diagram of $\mathcal{Q}_{\rm min}$ as a function of $\tilde{\Omega}$ and $K$ (Fig. 4) by simultaneously scanning $\delta\omega_{1}$ and $\mathcal{A}$.
@@
Note that $\langle j_{1}\rangle$ is dictated by the typical current-coherence relation (Eq. (22)), whereas $\langle j_{2}\rangle$ is associated only with the imaginary part of the correlated coherence, ${\rm Im}[\rho^{\rm ss}_{10,01}]$. It is noteworthy that another correlated coherence ${\rm Im}[\rho^{\rm ss}_{11,00}]$ (Fig. S1C) makes no contribution to the current. This is most likely associated with the fact that $\langle j_{2}\rangle>0$ (Fig. 2D) for all $\delta\omega_{1}$ and $\delta\omega_{2}$ is more consistent with ${\rm Im}[\rho^{\rm ss}_{10,01}]$ (Fig. 2C) than ${\rm Im}[\rho^{\rm ss}_{11,00}]$, the latter of which is not always positive (Fig. S1C). Despite the inherent asymmetry, the currents generated from the 1st and 2nd qubits become, on average, comparable when the two qubits are strongly coupled at large $K$ (data points in green, Fig. 5B).
@@ IV.3 TUR of subsystem and comparison of TQS with two coupled noisy oscillators
(i) For $K_{os}=0$, the two oscillators exhibit phase dynamics fully decoupled from each other [15]. As a result, the total entropy production, the variance and the mean phase angle for $\tau\gg 1$ from each oscillator ($i=1$ and 2) are given by
@@
Thus, the uncertainty product for the $i$th oscillator is
@@
To draw an analogy with the TQS where only the 1st qubit is irradiated by the external field, we set $\Omega_{1}$ finite and $\Omega_{2}\rightarrow 0$ for the two coupled oscillators. Two limiting cases of the oscillators and their comparisons with the TQS illuminate the fundamental differences between the couplings in the TQS and in the classical oscillators. (i) For $K_{os}\rightarrow 0$, the uncertainty product of the 1st oscillator is minimized to the TUR bound, whereas the 2nd oscillator with $\Omega_{2}\rightarrow 0$ is merely subject to the thermal noise, so that we get $\mathcal{Q}_{1}=2$ and $\mathcal{Q}_{2}\gg 1$. This trend is also observed in the TQS when $\tilde{\Omega}$ is finite and $K\rightarrow 0$. Figure 6A shows the uncertainty products of the total system ($\mathcal{Q}$) and the subsystems ($\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$) over the detunings. As previously discussed, the TQS effectively behaves as two non-interacting qubits. The 1st qubit subject to the external field acquires the minimal uncertainty product $\mathcal{Q}_{1}\approx 1.25$ at $\delta\omega_{1}\approx 0$. Meanwhile, as the 2nd qubit is effectively decoupled from the 1st qubit with no source of driving, its uncertainty product is large ($\mathcal{Q}_{2}\gg 1$, Fig. 6A). (ii) For $K_{os}\gg|\Omega_{1}-\Omega_{2}|\approx\Omega_{1}$, the two oscillators are fully synchronized. Then, the current, its fluctuations, and the entropy production are reduced to half the values, yielding the uncertainty product minimized to its bound, $\mathcal{Q}_{1}=\mathcal{Q}_{2}=2$. Interestingly, in the corresponding regime for TQS ($K\gg 1$), Fig 6B indicates that the uncertainty product of each qubit is significantly large ($\mathcal{Q}_{1,2}\gg 1$) particularly under off-resonance condition. Even under the resonance condition, each qubit’s uncertainty product always exceeds 2, while the uncertainty product for the total system is still around the minimal bound ($\mathcal{Q}\approx 1.25$) (see also Fig. S4 and Fig. 2D, which give rise to $\mathcal{Q}$ and $\mathcal{Q}_{1,2}$). Note that the correlated coherence, particularly the superposition of |$\ket{01}$ and |$\ket{10}$ basis states (${\rm Im}[\rho_{10,01}]$, see Fig. 2C) contributes to the current from each qubit (Eqs. (32) and  (33)), while the total current originates only from the coherences of the 1st qubit (Eq. (31)). This is a feature unique to the TQS, but lacking in the two coupled oscillators.
@@ IV.4 Effect of entanglement on the TUR bound
Any meaningful signature of entanglement is found for $K\geq 1$ and $\tilde{\Omega}\geq 1$ (Fig. 7). The parameter space giving rise to a strong entanglement does not necessarily coincide with the region with small $\mathcal{Q}$ (compare Fig. 3A and Fig. 7). As vividly shown by the map of the smallest eigenvalue of the partially transposed density matrix ($\Lambda(\delta\omega_{1},\delta\omega_{2})$), the entanglement between the two qubits characterized with $\Lambda<0$ is the most apparent at $\tilde{\Omega}=10$ and $K=10$; yet, it contributes little to loosening the TUR bound, a finding consistent with a recent study on a coherent mesoscopic transport along a double quantum dot coupled to two thermal reservoirs [40]. The overlap between the two maps simultaneously satisfying $\mathcal{Q}(\delta\omega_{1},\delta\omega_{2})\lesssim 2$ and $\Lambda(\delta\omega_{1},\delta\omega_{2})<0$, quantified using the overlap parameter $\chi=\Theta(2-\mathcal{Q})(2/\mathcal{Q})\Lambda$, is shaped along the resonance condition ($\delta\omega_{1}\delta\omega_{2}=K^{2}$) only for $\tilde{\Omega}=1$ and $K=1$, 10 (Fig. 7B).
IV.5 Efficiency of a stochastic engine made of TQS
Given the quantumness characterizing the TQS, which lowers the TUR bound, it is of great interest to consider using TQS as a stochastic engine to convert available energy into useful work. With an assumption that the laser driving power $P$ is transmitted to the TQS via excitation, the maximum amount of power that can be extracted from TQS is $\dot{w}_{\rm max}=\hbar\omega\langle j\rangle$, which defines the efficiency, $\eta$, of the TQS as an engine
$\eta=\frac{\dot{w}_{\rm max}}{P}.$ (43)
Since the laser power is consumed to perform work $\dot{w}_{\rm max}$, the heat current from the system to the surrounding bath is given as $T\Sigma^{ss}=P-\hbar\omega\langle j\rangle$. From the inequality of the uncertainty product that specifies the TUR bound, the lower bound of non-useful heat current is written as [4]
$T\Sigma^{\rm ss}\geq\mathcal{Q}_{\rm min}k_{B}T\langle j\rangle^{2}/\text{Var}(j).$ (44)
Thus, it follows that the efficiency for the TQS is upper-bounded as
$\eta\leq\eta_{\rm TUR}$ $=\frac{1}{\mathcal{Q}_{\rm min}k_{B}T\left(\frac{\langle j\rangle}{\hbar\omega\text{Var}(j)}\right)+1}=\frac{1}{\frac{\mathcal{Q}_{\rm min}}{\mathcal{Q}}+1},$ (45)
where the relation $\mathcal{Q}=\mathcal{A}\mathcal{F}$ in Eq. (16) with $\mathcal{A}=\Delta S_{\rm tot}/k_{B}\langle n\rangle=\hbar\omega/k_{B}T$ and $\mathcal{F}=\text{Var}(j)/\langle j\rangle$ is used to obtain the last equality. The expression for the maximum efficiency set by the TUR bound ($\mathcal{Q}_{\rm min}$) suggests that when the TQS is employed as an engine to extract work, its efficiency is upper-bounded by $\eta_{\rm TUR}$. For TQS, the TUR bound ($\mathcal{Q}_{\rm min}^{\rm TQS}$) is smaller than the bound of classical counterpart, i.e., $\mathcal{Q}_{\rm min}^{\rm TQS}<2$, suggesting that the quantumness present in the TQS can enhance the engine efficiency.
IV.6 Effect of quantum nature of external driving field
Although the current study adopts the semi-classical description of light source, from the quantum-mechanical perspective of light-matter interaction, it is still of interest to explore the effect of quantized light, particularly, on the TUR bound. In the regimes involving relatively strong external field, with Rabi frequency $\Omega$, while satisfying the requirement of RWA ($\omega\gg\Omega$), the field irradiating the TQS can be represented as a coherent state and its interaction with quantum objects is described using Jaynes-Cummings Hamiltonian,
$H_{\rm JC,int}=\hbar g(\hat{a}\sigma_{+}+\hat{a}^{\dagger}\sigma_{-}),$ (46)
where $g$ is a coupling strength, and $\hat{a}$ and $\hat{a}^{\dagger}$ are the creation and annihilation operators of the quantized light field. Here, a coherent state of light is described as |$\ket{\alpha}=e^{-|\alpha|^{2}/2}\sum_{n=0}^{\infty}\frac{\alpha^{n}}{\sqrt{n!}}\ket{n}$ expanded in occupation number representation, satisfying the relation $\hat{a}\ket{\alpha}=\alpha\ket{\alpha}$. We note that in the limit of strong field, i.e., $|\alpha|^{2}\gg 1$, coherent states also approximately satisfy the relation, $\bra{\alpha}e^{-(i/\hbar)H_{\rm JC,int}t}\ket{\alpha}\approx e^{-ig(\alpha\sigma_{+}+\alpha^{*}\sigma_{-})t}$, which reduces to the semi-classical description of the light-matter interaction in Eq. (5). Under this condition, our semi-classical interaction Hamiltonian remians valid, and the effect of quantized fields is only marginal. In the weak field regime, on the other hand, higher-order corrections are no longer insignificant, and are likely contribute to the nonequilibrium current and current fluctuation.
In addition, non-classical states of light, such as squeezed states, could significantly influence both quantum coherence and consequently current. This influence has a potential to lower the TUR bound and, in turn, increase the engine efficiency beyond the classical limit [62, 63, 64].
@@ V Summary
    Provided that the external field is strong enough ($\tilde{\Omega}\geq 1$), a strong coupling ($K\gg 1$) between the two qubits is essential for creating an entanglement in the TQS, and for transmitting excitation and coherence from the 1st qubit to the 2nd qubit (compare Fig. 2A, B with Fig. S1A, B). The 1st and 2nd qubits share the current generation with increasing coupling strength, while the overall photon current from the system is attenuated (Eqs. (31)– (33) and Fig. 2D).
@@
    The most effective suppression of photon current fluctuations, which gives rise to TUR violation, is achieved when $\tilde{\Omega}$ is moderate, and the system satisfies the proper detuning condition ($\delta\omega_{1}\delta\omega_{2}=K^{2}$), the latter of which underscores the aspect of light-matter interaction of the current study and distinguishes it from the recent two-qubit model considered in the context of a temperature gradient-driven electron-transport against chemical gradient [40].
@@
Here are our final remarks with prospect. The suboptimal bound of the uncertainty product, $\mathcal{Q}_{K\gg 1,\tilde{\Omega}>3,{\rm min}}^{\rm TQS}\approx 1.36$, which is acquired for strongly coupled qubits subject to a strong field (Fig. 4), is specific to the TQS. Nevertheless, together with our finding from Fig. 7 that the entanglement between two qubits plays little role in loosening the TUR bound, the inequality of Eq. (49) obtained based on TQS may be general enough to be extended to multi-qubit systems. We surmise that the TLS sets the fundamental limit of TUR even for multi-qubit systems. As an extension of TQS, the $N$-qubit system can be related to photosynthetic systems, more specifically light-harvesting complexes (LHC) comprised of quantum-mechanically coupled multiple chromophores irradiated by the light source [65, 66, 67, 68]. The problem of how the TUR of mutually or nearest-neighbor coupled qubits changes would be a topic of great interest for future investigation.
@@ VI.2 Method of cumulant generating function
where we have introduced the total emission and absorption operators from the two qubits, ±=±$\mathcal{L}_{\pm}=\mathcal{L}^{1}_{\pm}+\mathcal{L}^{2}_{\pm}$ and defined the remaining elements as $\mathcal{L}_{0}\equiv\mathcal{L}-\mathcal{L}_{+}-\mathcal{L}_{-}$. The two last terms on the right hand side of Eq. (54) express the contributions of emission and absorption to the evolution of $\tilde{\rho}(n,t)$ with $\mathcal{L}_{\pm}$. For the TQS, these operators can be written explicitly as:
@@
To obtain the current from each qubit, the total absorption and emission operators can be redefined as ±=±$\mathcal{L}_{\pm}=\mathcal{L}^{1}_{\pm}$ or ±$\mathcal{L}^{2}_{\pm}$ instead of summing up the entire contribution from the two qubits as ±=±$\mathcal{L}_{\pm}=\mathcal{L}^{1}_{\pm}+\mathcal{L}^{2}_{\pm}$. The expressions for the Liouville operators associated with emissions ($\mathcal{L}_{+}^{1}$, $\mathcal{L}_{+}^{2}$) and absorptions ($\mathcal{L}_{-}^{1}$, $\mathcal{L}_{-}^{2}$) given in Eq. (55) offer the expressions in Eqs. (29) and (30).
@@ Acknowledgements.
Supplemental Information
Each of the density matrix elements in Eq. (11) evolves as follows.
$\dot{\rho}_{11,11}$ $=-[\gamma(\bar{n}_{2}+1)+\left(\bar{n}_{1}+1\right)]\rho_{11,11}+\bar{n}_{1}\rho_{01,01}+\gamma\bar{n}_{2}\rho_{10,10}-i\underbrace{\left(-\Omega\rho_{01,11}+\Omega\rho_{11,01}\right)}_{=2i\Omega{\rm Im}[\rho_{11,01}]}$
$\dot{\rho}_{11,10}$ $=\bar{n}_{1}\rho_{01,00}-\left[(\bar{n}_{1}+1)+\frac{\gamma}{2}(2\bar{n}_{2}+1)\right]\rho_{11,10}-i\left(-\Omega\rho_{01,10}+\Omega\rho_{11,00}-K\rho_{11,01}+\delta\omega_{2}\rho_{11,10}\right)$
$\dot{\rho}_{11,01}$ $=\gamma\bar{n}_{2}\rho_{10,00}-\left[\frac{1}{2}(2\bar{n}_{1}+1)+\gamma(\bar{n}_{2}+1)\right]\rho_{11,01}-i\left(-\Omega\rho_{01,01}+\delta\omega_{1}\rho_{11,01}-K\rho_{11,10}+\Omega\rho_{11,11}\right)$
$\dot{\rho}_{11,00}$ $=-\left[\frac{1}{2}(2\bar{n}_{1}+1)+\frac{\gamma}{2}(2\bar{n}_{2}+1)\right]\rho_{11,00}-i\left(-\Omega\rho_{01,00}+(\delta\omega_{1}+\delta\omega_{2})\rho_{11,00}+\Omega\rho_{11,10}\right)$
$\dot{\rho}_{10,11}$ $=\bar{n}_{1}\rho_{00,01}-\left[(\bar{n}_{1}+1)+\frac{\gamma}{2}(2\bar{n}_{2}+1)\right]\rho_{10,11}-i\left(-\Omega\rho_{00,11}+K\rho_{01,11}+\Omega\rho_{10,01}-\delta\omega_{2}\rho_{10,11}\right)$
$\dot{\rho}_{10,10}$ $=\bar{n}_{1}\rho_{00,00}-\left[(\bar{n}_{1}+1)+\gamma\bar{n}_{2}\right]\rho_{10,10}+\gamma(\bar{n}_{2}+1)\rho_{11,11}-i\underbrace{\left[\Omega(\rho_{10,00}-\rho_{00,10})+K(\rho_{01,10}-\rho_{10,01})\right]}_{=2i\Omega{\rm Im}[\rho_{10,00}]-2iK{\rm Im}[\rho_{10,01}]}$
$\dot{\rho}_{10,01}$ $=-\left[\frac{1}{2}(2\bar{n}_{1}+1)\rho_{10,01}+\frac{\gamma}{2}(2\bar{n}_{2}+1)\right]\rho_{10,01}-i\left(-\Omega\rho_{00,01}+K\rho_{01,01}+(\delta\omega_{1}-\delta\omega_{2})\rho_{10,01}-K\rho_{10,10}+\Omega\rho_{10,11}\right)$
$\dot{\rho}_{10,00}$ $=-\left[\frac{1}{2}(2\bar{n}_{1}+1)+\gamma\bar{n}_{2}\right]\rho_{10,00}+\gamma(\bar{n}_{2}+1)\rho_{11,01}-i\left(-\Omega\rho_{00,00}+K\rho_{01,00}+\delta\omega_{1}\rho_{10,00}+\Omega\rho_{10,10}\right)$
$\dot{\rho}_{01,11}$ $=\gamma\bar{n}_{2}\rho_{00,10}-\left[\frac{1}{2}(2\bar{n}_{1}+1)+\gamma(\bar{n}_{2}+1)\right]\rho_{01,11}-i\left(\Omega\rho_{01,01}-\delta\omega_{1}\rho_{01,11}+K\rho_{10,11}-\Omega\rho_{11,11}\right)$
$\dot{\rho}_{01,10}$ $=-\left[\frac{1}{2}(2\bar{n}_{1}+1)+\frac{1}{2}\gamma(2\bar{n}_{2}+1)\right]\rho_{01,10}-i\left(\Omega\rho_{01,00}-K\rho_{01,01}-(\delta\omega_{1}-\delta\omega_{2})\rho_{01,10}+K\rho_{10,10}-\Omega\rho_{11,10}\right)$
$\dot{\rho}_{01,01}$ $=\gamma\bar{n}_{2}\rho_{00,00}-[\bar{n}_{1}+\gamma(\bar{n}_{2}+1)]\rho_{01,01}+(\bar{n}_{1}+1)\rho_{11,11}-i\underbrace{\left[K(\rho_{10,01}-\rho_{01,10})-\Omega(\rho_{11,01}-\rho_{01,11})\right]}_{=2iK{\rm Im}[\rho_{10,01}]-2i\Omega{\rm Im}[\rho_{11,01}]}$
$\dot{\rho}_{01,00}$ $=-\left[\bar{n}_{1}+\frac{\gamma}{2}(2\bar{n}_{2}+1)\right]\rho_{01,00}+(\bar{n}_{1}+1)\rho_{11,10}-i\left(\delta\omega_{2}\rho_{01,00}+K\rho_{10,00}+\Omega\rho_{01,10}-\Omega\rho_{11,00}\right)$
$\dot{\rho}_{00,11}$ $=-\left[\frac{1}{2}(2\bar{n}_{1}+1)+\frac{\gamma}{2}(2\bar{n}_{2}+1)\right]\rho_{00,11}-i\left(\Omega\rho_{00,01}-(\delta\omega_{1}+\delta\omega_{2})\rho_{00,11}-\Omega\rho_{10,11}\right)$
$\dot{\rho}_{00,10}$ $=-\left[\frac{1}{2}(2\bar{n}_{1}+1)+\gamma\bar{n}_{2}\right]\rho_{00,10}+\gamma(\bar{n}_{2}+1)\rho_{01,11}-i\left(\Omega\rho_{00,00}-K\rho_{00,01}-\delta\omega_{1}\rho_{00,10}-\Omega\rho_{10,10}\right)$
$\dot{\rho}_{00,01}$ $=-\left[\bar{n}_{1}+\frac{\gamma}{2}(2\bar{n}_{2}+1)\right]\rho_{00,01}+(\bar{n}_{1}+1)\rho_{10,11}-i\left(-\delta\omega_{2}\rho_{00,01}-K\rho_{00,10}-\Omega\rho_{10,01}+\Omega\rho_{00,11}\right)$
$\dot{\rho}_{00,00}$ $=-(\bar{n}_{1}+\gamma\bar{n}_{2})\rho_{00,00}+\gamma(\bar{n}_{2}+1)\rho_{01,01}+(\bar{n}_{1}+1)\rho_{10,10}-i\underbrace{\left(\Omega\rho_{00,10}-\Omega\rho_{10,00}\right)}_{=-2i\Omega{\rm Im}[\rho_{10,00}]}$ (S1)
The steady state conditions for the population, $\dot{\rho}_{11,11}=0$, $\dot{\rho}_{10,10}=0$, $\dot{\rho}_{01,01}=0$, and $\dot{\rho}_{00,00}=0$, yield
$2\Omega{\rm Im}[\rho_{11,01}^{\rm ss}]=[(\bar{n}_{1}+1)\rho_{11,11}^{\rm ss}-\bar{n}_{1}\rho_{01,01}^{\rm ss}]+[\gamma(\bar{n}_{2}+1)\rho_{11,11}^{\rm ss}-\gamma\bar{n}_{2}\rho_{01,01}^{\rm ss}]$ (S2)
$-2\Omega{\rm Im}[\rho_{10,00}^{\rm ss}]+2K{\rm Im}[\rho_{10,01}^{\rm ss}]=-[(\bar{n}_{1}+1)\rho_{10,10}^{\rm ss}-\bar{n}_{1}\rho_{00,00}^{\rm ss}]+[\gamma(\bar{n}_{2}+1)\rho_{11,11}^{\rm ss}-\gamma\bar{n}_{2}\rho_{10,10}^{\rm ss}]$ (S3)
$2\Omega{\rm Im}[\rho_{11,01}^{\rm ss}]-2K{\rm Im}[\rho_{10,01}^{\rm ss}]=[(\bar{n}_{1}+1)\rho_{11,11}^{\rm ss}-\bar{n}_{1}\rho_{01,01}^{\rm ss}]-[\gamma(\bar{n}_{2}+1)]\rho_{01,01}^{\rm ss}-\gamma\bar{n}_{2}\rho_{00,00}^{\rm ss}]$ (S4)
$2\Omega{\rm Im}[\rho_{10,00}^{\rm ss}]=[(\bar{n}_{1}+1)\rho_{10,10}^{\rm ss}-\bar{n}_{1}\rho_{00,00}^{\rm ss}]+[\gamma(\bar{n}_{2}+1)\rho_{01,01}^{\rm ss}-\gamma\bar{n}_{2}\rho_{00,00}^{\rm ss}]$ (S5)
Therefore, together with the current from each qubit given in Eqs. (29) and (30), subtracting Eq. (S4) from Eq. (S2) or adding Eq. (S3) and Eq. (S5) yields Eq. (33), whereas adding Eq. (S2) and Eq. (S5) yields Eq. (31).
Figure S1: Real and imaginary parts of coherence of the 2nd qubit when the 1st qubit is in its (A) excited and (B) ground states. (C) The correlated coherence between |$\ket{11}$ and |$\ket{00}$.
Figure S2: Photon current fluctuations over varying amount of detunings calculated for 9 different sets of parameters ($K=0.1$, 1, 10 and $\tilde{\Omega}=0.1$, 1, 10).
Figure S3: $\mathcal{Q}$ as a function of $K$ and $\mathcal{A}$ under the constraints of $\delta\omega_{1}=K\cdot\tilde{\Omega}$ and $\delta\omega_{2}=K/\tilde{\Omega}$. We also made another constraint $\tilde{\Omega}$. For large $K(\gg 1)$, and thus large $\tilde{\Omega}$, the minimal $\mathcal{Q}$ is obtained when $\mathcal{A}\approx 4.08$.
Figure S4: Photon current fluctuations of entire system and each qubit as a function of detunings ($\delta\omega_{1},\delta\omega_{2}$).